OpenStudy (anonymous):

121° 22' 36" is the same as _____°.

OpenStudy (anonymous):

use, 1' = 1/60 degree and , 1'' = 1/60 minute so , 1'' = 1/3600 degree

OpenStudy (anonymous):

121 degree + 22' = 22/60 degree + 36'' = 36/3600 degree = ?????

OpenStudy (anonymous):

121.376?

OpenStudy (unklerhaukus):

6 rounds up, not down

OpenStudy (anonymous):

yeah I mean 121.38 But that's the right answer, right?

OpenStudy (unklerhaukus):

yeah, that is an approximation of the exact value, if you want to avoid error leave it as a fraction